# Safe On-Orbit Dislodging of Deployable Structures via Robust Adaptive MPC

### v1

@@ IV Simulation results
Figure 4: Dislodging mission using a servicer which contains simulated on MuJoCo Platform in a zero-gravity environment.
@@
In this section, we formulate the discrete-time linear time-varying system of the client and the performance of our robust adaptive MPC is also presented in this section with the comparison with a baseline adaptive MPC algorithm in [25]. We choose the MuJoCo [32] as our simulation platform to emulate both client and servicer in a zero-gravity environment. The parameters about the client are defined in Table. I.
The uncertainty in the parameters is described by $\mathcal{P}=\left\{\boldsymbol{\rho}_{k}\in\mathbb{R}^{3}\bigr{|}\>||\boldsymbol{\rho}_{k}||_{\infty}\leq 8\right\},$ with $\boldsymbol{\rho}^{*}_{k}=[\mathcal{K}_{k},\zeta_{k},\mathcal{D}_{k}]^{\top}$. The disturbance set is $\mathbb{W}=\left\{w\in\mathbb{R}^{3}\bigr{|}\>||w||_{\infty}\leq 0.2\right\}$ and the state and input constraints are described by
$\mathbb{Z}$ =$=\left\{(\mathbf{x},\mathbf{u}){\in}\mathbb{R}^{2\times 6}\left|\>\begin{array}[]{rl}||\mathbf{x}[1]||\leq 0.7&,||\mathbf{x}[2]||\leq 0.2\\ \left\lVert\mathbf{u}[1,..,3]\right\rVert\leq 50&,\left\lVert\mathbf{u}[4,..,6]\right\rVert\leq 10\\ \end{array}\right.\right\}.$
The initial state of the system is $\mathbf{x}_{0}=[0.6,0.2]^{\top}$. In both AMPC and our method, the state tube is constructed by translating and scaling the set $\mathbb{X}_{0}=\left\{\mathbf{x}\in\mathbb{R}^{2}\bigr{|}\>\left\lVert\mathbf{x}[1]\right\rVert\leq 0.8,\left\lVert\mathbf{x}[2]\right\rVert\leq 0.25\right\}$. The bounded complexity update of $\mathcal{P}_{k}$ is performed using $n_{\theta}=45$ hyperplanes which are initially chosen as outer bounds of the set $\mathcal{P}$. The cost matrices are given as
$\begin{array}[]{l l l}\mathbf{Q}=\begin{bmatrix}1.47&0\\ 0&1.35\end{bmatrix},&\>\mathbf{R}=\begin{bmatrix}0.94&0\\ 0&1.0\end{bmatrix},&\>\boldsymbol{\Upsilon}=\begin{bmatrix}1.15&0\\ 0&1.25\end{bmatrix}.\end{array}$
@@
[…]
@@
Fig. 6(a) to Fig. 6(c) shows the results of estimating the parameter within polytope based on our method compared with two baseline methods. All parameter membership sets reduce in polygon size during the estimation process. Compare Fig. 6(a) for adaptive control based on our previous work and Fig. 6(b) for AMPC, our method results in Fig. 6(c) shows faster speed to catch the true parameter value than baseline methods. Note that the adaptation of parameters is influenced by both the initial conditions and the realization of disturbances, as the cost function does not inherently account for the benefits of future parameter learning.
Fig. 6(d) to Fig. 6(f) shows the results of each time-varying parameter estimation during the dislodging. The shadow area shows the bounded values for the estimates in each time step. Based on the result of our method compared with the two baseline methods, our method shows that it can get the true value faster, and the error is much smaller than the two baseline methods.
@@ V Experiment Results
To prove the feasibility of our robust adaptive MPC in hardware, we also implement our robust adaptive MPC in our testbed compared with PID and adaptive control, as shown in Fig. 7. We leverage the testbed in our lab to build up the dislodging environment in space but consider the gravity based on the limitation. We use UR5e to emulate a Servicer that can apply 6 DoF wrench in SE(3). Also, we use the WAM robot arm to hold the satellite module, the details of which are shown on the right-bottom side, to emulate a Client, which is a free-flying spacecraft that can decouple the dynamics from the Servicer during dislodging. For both Client and Servicer, we set up two types of safety constraints in our algorithm as shown in Table. II based on the hardware limitation to ensure the dislodging can proceed well as shown in Fig. 8a; once either is out of its hardware constraints based on the value in Table. II, the system will stop to denote the potential risk occurrence during the dislodging as shown in Fig. 8b. Note that to compare and discover the performance with different methods, we set up a scheme in our program that if any variable’s value shown in Table II is between the hardware constraints and software constraints, the robot system can still work but denote the potential damage will happen to simulate the real scenarios in space. The hinge’s top-side view during dislodging is shown in Fig. 9. All the safe constraints we set up in our program are under the real constraints for both Client and Servicer to avoid any danger happening on the hardware.
@@
Fig. 10 shows the wrench applied by the Client during the dislodging using PID control method. Based on the results, we can see clearly that the PID method can make the Client out of its safe range easily, which will break the hardware easily.
@@
Fig. 11 shows the wrench applied by the Client during the dislodging via adaptive control. Based on the results, it’s also straightforward to see that even though the adaptive control method shows a little bit better than PID, it still can make the Client be out of its safe range to bring the potential breakage for the hardware.
@@
Fig. 12 shows the wrench applied by the Client during the dislodging via our robust adaptive MPC. Based on the results, we can find that the adaptive control method can ensure the wrench from the Client is always under the constraint to keep the safe during dislodging.

### v2

@@ IV Simulation results
Figure 4: Dislodging mission using a servicer which contains simulated on MuJoCo Platform in a freefall environment.
@@
In this section, we formulate the discrete-time linear time-varying system of the client and the performance of our robust adaptive MPC is also presented in this section with the comparison with a baseline adaptive MPC algorithm in [aboudonia2024adaptive]. We choose the MuJoCo [todorov2012mujoco], an advanced robotics physical platform which can fully emulate real space environment, as our simulation platform to emulate both client and servicer in a freefall environment. The parameters about the client are defined in Table. I. Note that all parameters we selected in this section are to be physically reasonable and representative of our experimental hardware setup.
[…]
The initial state of the Newton system is $\mathbf{x}_{0}=[0.6,0.2]^{\top}$. We set up the sampling period parameter $h$ in Eq. 6 as $0.01s$ to maintain consistency for on the hardware feature and simulation platform. In both AMPC and our method, the state tube is constructed by translating and scaling the set $\mathbb{X}_{0}=\left\{\mathbf{x}\in\mathbb{R}^{2}\bigr|\>\left\lVert\mathbf{x}[1]\right\rVert\leq 0.8,\left\lVert\mathbf{x}[2]\right\rVert\leq 0.25\right\}$. The bounded complexity update of $\mathcal{P}_{k}$ is performed using $n_{\theta}=45$ hyperplanes which are initially chosen as outer bounds of the set $\mathcal{P}$. The cost matrices are given as
[…]
@@
Experimental Value 1.57 rad -1.57 rad 0.5 rad/s -0.5 rad/s $\cdotm -10 N\cdotm$ 50 N -50 N
Simulation Value 0.3 rad -1.2 rad 0.2 rad/s -0.2 rad/s $\cdotm -2 N\cdotm$ 20 N -20 N
@@
Fig. 6 presents a detailed comparison of the exploration performance, illustrating the evolution of the time-varying parameter estimation for our robust adaptive MPC (RAMPC) against the baseline adaptive control (AC) and AMPC methods. Fig. 6(a) to Fig. 6(c) depict the 3D evolution of the parameter membership set (polytope) over a 20s horizon. While all three controllers successfully reduce the size of the uncertainty polytope from its initial bounds, a stark difference in convergence efficacy is evident. The baseline AC in Fig. 6 and AMPC in Fig. 6(b) controllers, which rely primarily on passive excitation, demonstrate a slow and less significant reduction, resulting in a large final uncertainty set. In striking contrast, our RAMPC in Fig. 6(c) achieves a significantly more rapid and aggressive tightening of the parameter set, converging to a much smaller final polytope that closely encloses the true parameter trajectory, visualized as the dark inner line. This superior performance is a direct consequence of the novel dual-mode cost in (25), which explicitly penalizes the size of the predicted parameter uncertainty set. This term actively guides the controller to balance exploration to excite the system and reduce uncertainty with exploitation, thereby enhancing control performance. which is the feature that the baseline methods lack.
This quantitative advantage is further clarified in Fig. 6(d) to Fig. 6(f), which plot the bounded estimates for each individual time-varying parameter $\mathcal{K}$, $\zeta$, and $\mathcal{D}$ against time. The shaded regions represent the bounds of the estimated parameter set at each time step, while the dashed line indicates the ground-truth, time-varying parameter value. Across all three plots, the bounds generated by our RAMPC converge more rapidly and remain significantly tighter around the true value compared to both the AMPC and AC baselines. This visualization confirms that our method not only achieves a faster rate of learning but also maintains a much smaller estimation error throughout the dislodging process. This robust parameter identification is critical, as it allows the controller to operate with less conservative bounds, directly improving control performance and safety, as validated in the state-tracking results shown in Fig. 5.
@@ V Experiment Results
To prove the feasibility of our robust adaptive MPC in hardware, we also implement our robust adaptive MPC in our testbed compared with PID and adaptive control, as shown in Fig. 7. We leverage the testbed in our lab to simulate the space dislodging environment considering the $1g$ gravity based on the terrestrial laboratory. We use UR5e [^3] https://www.universal-robots.com/products/ur5e/, by Universal Robot. to emulate a Servicer that can apply 6 DoF wrench in SE(3). Also, we use the WAM robot arm [^4] https://barrett.com/wam, by Barrett Technology. to hold the satellite module, the details of which are shown on the right-bottom side, to emulate a Client, which is a free-flying spacecraft that can decouple the dynamics from the Servicer during dislodging. For both Client and Servicer, we set up two types of safety constraints in our algorithm as shown in Table. II based on the hardware limitation to ensure the dislodging can proceed well as shown in Fig. 10a; once either is out of its hardware constraints based on the value in Table. II, the system will stop to denote the potential risk occurrence during the dislodging as shown in Fig. 10b. Note that to compare and discover the performance with different methods, we set up a scheme in our program that if any variable’s value shown in Table II is between the hardware constraints and software constraints, the robot system can still work but denote the potential damage will happen to simulate the real scenarios in space. The hinge’s top-side view during dislodging is shown in Fig. 8. All the safe constraints we set up in our program are under the real constraints for both Client and Servicer to avoid any danger happening on the hardware.
Table III presents a quantitative evaluation of the proposed RAMPC against baseline controllers, providing compelling evidence of its superior performance across both simulation and hardware implementations. The metrics clearly demonstrate that the RAMPC framework achieves the lowest root mean square errors for both angle and velocity tracking, signifying enhanced control accuracy and robustness. Furthermore, our method exhibits a substantially lower constraint violation rate, a critical outcome that validates its capacity for ensuring operational safety during the complex dislodging maneuver. The controller also attains the fastest convergence times in both experimental settings, highlighting its superior efficiency. These empirical results collectively underscore the primary contribution of this work, affirming that the proposed robust adaptive MPC strategy effectively delivers a more accurate, safe, and efficient solution for the on orbit servicing task under significant parametric uncertainty.
Fig. 9 presents a comparative analysis of the computational performance for the proposed Robust Adaptive Model Predictive Control RAMPC, a baseline Adaptive Model Predictive Control AMPC, and a Model Reference Adaptive Control MRAC scheme. The figure illustrates the runtime distribution in milliseconds for each control step, evaluated with a prediction horizon of six and employing warm starts to initialize the optimization. Both the RAMPC and AMPC methods exhibit comparable median computation times, which are observed to be well under two milliseconds. The interquartile ranges indicate consistent performance for both predictive controllers. In contrast, the non-predictive MRAC demonstrates substantially lower computational overhead, yet all three controllers comfortably satisfy the ten-millisecond control loop budget, confirming their suitability for real-time implementation.
@@
Fig. 11 shows the wrench applied by the Client during the dislodging using PID control method. Based on the results, we can clearly see that the PID method can easily cause the Client to exceed its safe range, which can damage the hardware.
@@
Fig. 12 shows the wrench applied by the Client during the dislodging via adaptive control. Based on the results, it’s also straightforward to see that even though the adaptive control method shows a little bit better than PID, it still can make the Client be out of its safe range, bring potential damage to the hardware.
Table III: Quantitative metrics from the plotted results under time varying uncertainty. Simulation uses zero gravity in MuJoCo with horizon six. Hardware uses the UR5e and WAM setup with software limits from Table II. Values are averages across trials estimated from the published figures.
Simulation Hardware
Metric RAMPC AMPC Adaptive Metric RAMPC Adaptive PID
Angle RMSE [rad] 0.003 0.004 0.005 Angle RMSE [rad] 0.320 0.380 0.430
Velocity RMSE [rad per second] 0.003 0.004 0.004 Velocity RMSE [rad per second] 0.080 0.100 0.120
Violation rate [per cent of steps] 0.04 0.5 1.5 Violation rate [per cent of steps] 0.2 1.0 3.0
Convergence time [s] 12.0 14.0 18.0 Convergence time [s] 12.0 18.0 22.0
@@
Fig. 13 shows the wrench applied by the Client during the dislodging via our robust adaptive MPC. Based on the results, we can see that the adaptive control method ensures the wrench from the Client is always within the constraint to maintain safety during dislodging.
@@
Figure 9: Runtime distributions in each step for RAMPC, AMPC, and MRAC at horizon six with warm starts. Boxes indicate the interquartile range and the center line denotes the median. Whiskers mark the lower and upper quantiles used for plotting and outliers are omitted. The vertical axis reports milliseconds per step. The control loop budget equals 10 ms.